# MAN TruckScenes: A multimodal dataset for autonomous trucking in diverse conditions

### v1

@@ Abstract
Autonomous trucking is a promising technology that can greatly impact modern logistics and the environment. Ensuring its safety on public roads is one of the main duties that requires an accurate perception of the environment. To achieve this, machine learning methods rely on large datasets, but to this day, no such datasets are available for autonomous trucks. In this work, we present MAN TruckScenes, the first multimodal dataset for autonomous trucking. MAN TruckScenes allows the research community to come into contact with truck-specific challenges, such as trailer occlusions, novel sensor perspectives, and terminal environments for the first time. It comprises more than 740 scenes of 20 s each within a multitude of different environmental conditions. The sensor set includes 4 cameras, 6 lidar, 6 radar sensors, 2 IMUs, and a high-precision GNSS. The dataset’s 3D bounding boxes were manually annotated and carefully reviewed to achieve a high quality standard. Bounding boxes are available for 27 object classes, 15 attributes, and a range of more than 230 m. The scenes are tagged according to 34 distinct scene tags, and all objects are tracked throughout the scene to promote a wide range of applications. Additionally, MAN TruckScenes is the first dataset to provide 4D radar data with 360° coverage and is thereby the largest radar dataset with annotated 3D bounding boxes. Finally, we provide extensive dataset analysis and baseline results. The dataset, development kit and more are available online.
@@ 1 Introduction
While numerous datasets for autonomous passenger cars exist [11], datasets for autonomous trucks are missing. However, heavy-duty vehicles have their own unique challenges. Large vehicles require different sensor mounting positions and rely on multiple sensors to cover the entire surrounding area. Moreover, trucks have to contend with occlusions from their own vehicle that change dynamically due to a movable truck-trailer combination and are affected by relative movements between their chassis and cabin. Furthermore, long-haul trucks operate in inherently different surroundings, such as logistics or container terminals. Therefore, a dedicated truck dataset is needed to develop reliable perception solutions for self-driving trucks.
@@ 2 Related Work
| KITTI [7] | nuScenes [3] | Waymo [15] | Argoverse2 [16] | ZOD [1] | Ours | |
| Scenes | 22 | 1000 | 1150 | 1000 | 1473 | 747 |
| Sample | ||||||
| Duration | ||||||
| Coverage | - | |||||
| Camera | 4 | 6 | 5 | 9 | 1 | 4 |
| Lidar | 1 | 1 | 5 | 2 | 3 | 6 |
| Radar | 0 | 5 | 0 | 0 | 0 | 6 |
| GNSS | 1 | 1 | 0 | 0 | 2 | 1 |
| IMU | 1 | 1 | 0 | 0 | 2 | 2 |
| Map | no | yes | no | yes | no | no |
| Range | ||||||
| Classes | 3 | 23 | 4 | 30 | 15 | 27 |
| Vehicle | car | car | car | car | car | truck |
@@
While there are numerous datasets focusing on various different aspects [11], all of the aforementioned datasets are limited to passenger cars, as shown in Table 1. To the best of our knowledge, there are only two datasets that include truck data. One is the TuSimple [18] dataset, which consists of 6408 images from a single front camera and provides annotations for lane markings only. The other is the SurMine [14] dataset, which is a proprietary dataset that only includes 2D bounding box annotations for 10665 camera images recorded in a mining facility. Hence, there are no large-scale perception datasets for autonomous trucking.
@@ 5 Experiments
We provide baseline 3D object detection results for the CenterPoint [17] detection model, which are shown in Table 3. The model is trained on fused ego-motion compensated lidar point clouds with 3 aggregated sweeps cropped to a grid with a voxel resolution of . The method uses seven detection heads and is evaluated for 12 detection classes, according to our metric definition. The results suggest that the detection quality of diverse object classes (like ’animal’ or ’other vehicle’) is insufficient and that the overall detection quality decreases significantly with increasing distance. Moreover, the results show a reduction in the detection quality under rain and fog conditions as well as a reduction within tunnels. In general, a more robust long-range perception is required for safe autonomous trucking.
| Method | Dataset | Modality | NDS | mAP | ATE | ASE | AOE | AVE | AAE |
| CenterPoint [17] | MAN | lidar | 0.43 | 0.30 | 0.40 | 0.35 | 0.29 | 2.93 | 0.20 |
@@ 6 Conclusion
In this work, we presented the first dataset for autonomous trucking. The MAN TruckScenes dataset is a large-scale multimodal dataset for the development of perception applications for autonomous trucks. The dataset contains a diverse set of scenes recorded in a multitude of different environmental conditions, provides coverage of all sensor modalities, and even 4D radar data. It includes carefully reviewed 3D bounding boxes, tracking information, and rich annotations for all objects within more than range. Furthermore, we introduced a generic method for compiling meaningful and balanced data splits. We provide baseline results for 3D object detection and promote research in all areas of perception, tracking, and prediction. For this purpose, we will maintain a public leaderboard and provide annotation instructions, taxonomy specifications, and a development kit. With this dataset, we aim to accelerate the development of autonomous trucking.
@@ A.6 Experiments
Experiments were conducted using a CenterPoint [17] model with a voxel-based encoder based on the SECOND architecture, a Feature Pyramid Network (FPN) neck, a Deep & Cross Network (DCN) CenterHead, and circular Non-maximum Suppression (NMS). The training scheme uses different data augmentation methods, including random global rotation, translation, scaling, flipping, point shuffling, and object sampling. The overall model is based on the MMDetection3D framework and was trained for 20 epochs on two Nvidia A100 GPUs for with a batch size of 16 per GPU.
@@
The baseline results of the CenterPoint model are shown in Table 3, while Table A1 shows the results per detection class. Observations show that the detection quality of diverse object classes (like ’animal’, ’other vehicle’, or ’bus’) is insufficient, while more homogeneous object classes (like ’traffic sign’, ’traffic cone’, or ’pedestrian’) have a higher detection score. Moreover, the size of the objects or the number of objects in the dataset are no sufficient predictors for the detection quality, as shown by the ’truck’ class.
@@
| car | 0.40 | 0.29 | 0.15 | 0.07 | 3.97 | 0.17 |
| truck | 0.26 | 0.41 | 0.16 | 0.03 | 2.65 | 0.39 |
| bus | 0.09 | 0.46 | 0.11 | 0.26 | 5.30 | 0.01 |
| trailer | 0.41 | 0.42 | 0.15 | 0.03 | 4.11 | 0.34 |
| other vehicle | 0.14 | 0.63 | 0.34 | 0.18 | 0.99 | 0.31 |
| pedestrian | 0.47 | 0.26 | 0.39 | 0.81 | 0.44 | 0.36 |
| motorcycle | 0.24 | 0.33 | 0.22 | 0.08 | 5.79 | 0.03 |
| bicycle | 0.05 | 0.27 | 0.27 | 0.46 | 2.12 | 0.14 |
| traffic cone | 0.59 | 0.28 | 0.39 | - | - | - |
| barrier | 0.43 | 0.20 | 0.60 | 0.04 | - | - |
@@
| traffic sign | 0.58 | 0.29 | 0.42 | 0.21 | - | 0.02 |
In addition, Table A2 provides an analysis of the detection quality over different range bins. The results suggest that the detection quality decreases with increasing distance. However, safe autonomous trucking requires a reliable long-range perception, especially for highway applications.
| CenterPoint | 0.53 / 0.54 | 0.46 / 0.51 | 0.34 / 0.45 | 0.30 / 0.43 |
Finally, MAN TruckScenes allows for the analysis of the detection quality by scene tags, which provides more insights into the model performance under different conditions. The results listed in Table A3 show the model performance under different weather, daytime, and lighting conditions, as well as the detection results for different areas, structures, and construction sides. The results suggest that rainy and foggy weather conditions negatively affect the model’s performance and that the model performs significantly worse in tunnels. However, the safe operation of autonomous trucks requires robust perception under diverse conditions.
| CenterPoint | ||
| weather | clear | 0.31 / 0.42 |
| overcast | 0.25 / 0.39 | |
| fog | 0.15 / 0.21 | |
| other_weather | - / - | |
| rain | 0.16 / 0.23 | |
| snow | - / - | |
| hail | - / - | |
| area | terminal | 0.33 / 0.33 |
| parking | - / - | |
| highway | 0.22 / 0.32 | |
| city | 0.18 / 0.29 | |
| residential | 0.18 / 0.31 | |
| rural | 0.18 / 0.29 | |
| other_area | - / - | |
| daytime | noon | 0.31 / 0.43 |
| morning | 0.27 / 0.37 | |
| night | - / - | |
| evening | 0.18 / 0.25 | |
| season | spring | - / - |
| summer | 0.39 / 0.44 | |
| autumn | 0.20 / 0.35 | |
| winter | 0.20 / 0.36 | |
| lighting | illuminated | 0.31 / 0.43 |
| twilight | 0.19 / 0.27 | |
| dark | 0.18 / 0.23 | |
| glare | 0.15 / 0.23 | |
| other_lighting | - / - | |
| structure | regular | 0.32 / 0.43 |
| underpass | 0.22 / 0.33 | |
| bridge | 0.21 / 0.37 | |
| overpass | 0.25 / 0.32 | |
| tunnel | 0.12 / 0.16 | |
| const | unchanged | 0.30 / 0.42 |
| roadworks | 0.25 / 0.30 |

### v2

@@ Abstract
Autonomous trucking is a promising technology that can greatly impact modern logistics and the environment. Ensuring its safety on public roads is one of the main duties that requires an accurate perception of the environment. To achieve this, machine learning methods rely on large datasets, but to this day, no such datasets are available for autonomous trucks. In this work, we present MAN TruckScenes, the first multimodal dataset for autonomous trucking. MAN TruckScenes allows the research community to come into contact with truck-specific challenges, such as trailer occlusions, novel sensor perspectives, and terminal environments for the first time. It comprises more than 740 scenes of 20 s each within a multitude of different environmental conditions. The sensor set includes 4 cameras, 6 lidar, 6 radar sensors, 2 IMUs, and a high-precision GNSS. The dataset’s 3D bounding boxes were manually annotated and carefully reviewed to achieve a high quality standard. Bounding boxes are available for 27 object classes, 15 attributes, and a range of more than 230 m. The scenes are tagged according to 34 distinct scene tags, and all objects are tracked throughout the scene to promote a wide range of applications. Additionally, MAN TruckScenes is the first dataset to provide 4D radar data with 360° coverage and is thereby the largest radar dataset with annotated 3D bounding boxes. Finally, we provide extensive dataset analysis and baseline results. The dataset, development kit, and more are available online.
@@ 1 Introduction
While numerous datasets exist for autonomous passenger cars [15], datasets for autonomous trucks are missing. However, heavy-duty vehicles have their unique challenges. Large vehicles require different sensor mounting positions and rely on multiple sensors to cover the entire surrounding area. Moreover, trucks have to contend with occlusions from their own vehicle that change dynamically due to a movable truck-trailer combination and are affected by relative movements between their chassis and cabin. Furthermore, long-haul trucks operate in inherently different surroundings, such as logistics or container terminals and their functionality has to be ensured under all environmental conditions to maintain a functioning logistics system. Therefore, a dedicated truck dataset is needed to develop reliable perception solutions for self-driving trucks under all conditions.
@@ 2 Related Work
Scenes | 22 | 1000 | 1150 | 1000 | 1473 | 176 | 21 | 747 | |||||||||||||||
Sample | |||||||||||||||||||||||
Duration | |||||||||||||||||||||||
Coverage | - | ||||||||||||||||||||||
Camera | 4 | 6 | 5 | 9 | 1 | 4 | 1 | 4 | |||||||||||||||
Lidar | 1 | 1 | 5 | 2 | 3 | 1 | 1 | 6 | |||||||||||||||
Radar | 0 | 5 | 0 | 0 | 0 | 2 | 1 | 6 | |||||||||||||||
GNSS | 1 | 1 | 0 | 0 | 2 | 1 | 1 | 1 | |||||||||||||||
IMU | 1 | 1 | 0 | 0 | 2 | 1 | 0 | 2 | |||||||||||||||
Map | no | yes | no | yes | no | no | no | no | |||||||||||||||
Range | |||||||||||||||||||||||
Classes | 3 | 23 | 4 | 30 | 15 | 14 | 13 | 27 | |||||||||||||||
Vehicle | car | car | car | car | car | car | car | truck |
@@
The aiMotive [17] dataset focuses on the collection of diverse driving scenes under severe weather conditions. It provides long-range annotations and data from a multimodal sensor setup for three distinct areas. However, the dataset’s extent is limited to a duration of and the two radar sensors do not cover a field of view (FoV).
The View-of-Delft [18] dataset is one of the first large-scale datasets with annotated 4D radar data and promotes the development of radar-based perception methods for vulnerable road users (VRUs) in city environments. Nevertheless, the geographical coverage of the dataset is limited, it does not include severe weather conditions, and only provides data from a single front-facing radar sensor.
While there are numerous other datasets focusing on various different aspects [15], all of the aforementioned datasets are limited to passenger cars, as shown in Table 1. To the best of our knowledge, there are only two datasets that include truck data. One is the TuSimple [25] dataset, which consists of 6408 images from a single front camera and provides annotations for lane markings only. The other is the SurMine [21] dataset, which is a proprietary dataset that only includes 2D bounding box annotations for 10665 camera images recorded in a mining facility. Hence, there are no large-scale or multimodal perception datasets for autonomous trucking.
@@ 5 Experiments
We provide baseline 3D object detection results for all three sensor modalities, which are shown in Table 3. The camera results are based on the PETR [16] model architecture with a pre-trained FCOS3D (V-99-eSE) backbone and an adjusted detection head to support 12-class classification. The detection distance is set to and the final image resolution is chosen to be pixels. The radar baseline is set by RadarGNN [10], which was trained on fused, ego-motion compensated 4D radar point clouds with 6 aggregated sweeps. The node features consisted of the rcs values, timestamps, connectivity degrees, and the relative radial velocities of the radar points. Lastly, a CenterPoint [24] model was trained on fused ego-motion compensated lidar point clouds with 3 aggregated sweeps cropped to a grid with a voxel resolution of . The model used seven detection heads and was evaluated for 12 detection classes, according to our metric definition.
The results suggest that the detection quality of diverse object classes (like ’animal’ or ’other vehicle’) is insufficient and that the overall detection quality decreases significantly with increasing distance. Moreover, the results show a reduced detection quality in winter conditions and tunnels. Furthermore, it can be observed that the camera model provides insufficient results for minority classes and struggles with challenging weather and lighting conditions, whereas the radar-based detection model provides insufficient results for all non-metallic object classes. Moreover, it can be observed that the radar performance is negatively affected by reflections within tunnels or container terminals. The lidar performance, on the other hand, is reduced under rainy or foggy conditions but also struggles within tunnel scenarios and seems to be more sensitive to the number of available training samples. In general, the results show that a more robust long-range perception is required for safe autonomous trucking.
Method | Dataset | Modality | mAP | NDS | ATE | ASE | AOE | AVE | AAE |
PETR [16] | MAN | camera | 0.02 | 0.12 | 1.13 | 0.69 | 0.65 | 1.50 | 0.56 |
RadarGNN [10] | MAN | radar | 0.07 | 0.11 | 0.89 | 0.81 | 1.13 | 8.00 | 0.57 |
CenterPoint [24] | MAN | lidar | 0.27 | 0.41 | 0.41 | 0.35 | 0.28 | 2.73 | 0.20 |
@@ 6 Conclusion
In this work, we presented MAN TruckScenes, the first dataset for autonomous trucking. It is a large-scale multimodal dataset for the development of autonomous truck perception applications. The dataset consists a diverse set of scenes recorded in a multitude of different environmental conditions and provides coverage of all sensor modalities, including 4D radar data. It contains carefully reviewed 3D bounding boxes, tracking information, and rich annotations for all objects within more than range. Furthermore, we introduced a generic method for compiling meaningful and balanced data splits. We provide baseline results for 3D object detection and promote research in all areas of perception, tracking, and prediction. For this purpose, we will maintain a public leaderboard and provide annotation instructions, taxonomy specifications, and a development kit. With this dataset, we aim to accelerate the development of autonomous trucking.
@@ 7 Limitations
8 Societal Impact
Despite the potential benefits of autonomous trucks the societal impact of the technology must be considered. Since autonomous trucks operate on public roads, their safety is a major concern that must be assured and cannot be compromised for economic reasons. While safety assurance is not just the key deployment factor, it is also the main driver for public acceptance and trust [7]. Therefore, regulatory authorities have to establish standards that ensure the safety of autonomous trucking and allow cross-border operation. Besides that, the impact on labor must be taken into account. While some studies expect job losses through autonomous vehicles [7], others argue that trained personnel will still be needed in the overall logistics process to supervise the system or carry out manual tasks [9]. Recent studies show that advanced automation can even help to improve working conditions and therefore, counteract the shortage of drivers and minimize turnover rates [8]. Ultimately, regulatory measures must be put in place to ensure that the interests of the general public are protected.
@@ A.6 Experiments
Experiments were conducted for all three sensor modalities. The camera results are based on the PETR [16] architecture, the radar baseline on RadarGNN [10], and the lidar results on CenterPoint [24]. To be specific, the camera baseline utilizes a PETR [16] model with a pre-trained FCOS3D (V-99-eSE) backbone with two multi-scale feature map outputs. Subsequently, a Feature Pyramid Network (FPN) neck is used for feature alignment and a transformer-based detection head is used with six decoder layers. The 900 query points are initialized within an area of and both the features as well as queries are positional encoded with a sine encoding. The training is Non Maximum Suppression (NMS) free and uses a Hungarian Assignment method instead, combined with a Focal Loss. The implementation is based on the MMDetection3D framework and uses common image augmentation techniques like random rotation, scaling, or translation. The final model is trained on an Nvidia A100 GPU for 48 epochs () and a batch size of 8.
The radar model uses the translation invariant RadarGNN [10] variant with a KNN-based graph construction method with a maximum of 20 neighbors in a radius of and directed edges with associated features. The main body of the model consists of five consecutive Message-Passing Neural Network (MPNN) layers and a separate classification and regression head. The training is based on a Cross-Entropy and Huber Loss function and does not utilize any data augmentation methods. The final training is conducted on an Nvidia A100 GPU with a batch size of 32 for 30 epochs.
@@
The lidar method is based on a CenterPoint [24] model with a voxel-based encoder based on the SECOND architecture, a FPN neck, a Deep & Cross Network (DCN) CenterHead, and circular NMS. The training scheme uses different data augmentation methods, including random global rotation, translation, scaling, flipping, point shuffling, and object sampling. The overall model is based on the MMDetection3D framework and was trained for 20 epochs on two Nvidia A100 GPUs for with a batch size of 16 per GPU.
In the following, detailed results of the different models are provided for every sensor modality. Table A1 - A3 show the class-specific results per sensor modality, while Table A4 lists the object detection results over different range bins. Table A5 shows the detection results by scene tags and Figure A12 - A14 provides qualitative results under various environmental conditions.
Observations show that the camera model struggles to perform well on minority classes and performs generally better on larger than smaller objects. The results show lots of false positives (FP) and the accumulation of bounding boxes on drivable areas. Besides that, the camera model’s performance decreases with increasing distance and performs significantly worse in the dark or tunnel environments.
The radar model, on the other hand, struggles to detect non-metallic objects but shows promising results on the majority classes. The results of the radar experiments also suggest that the accurate prediction of object orientations is most challenging and the velocity estimation seems to be off due to the utilization of the relative radial velocity instead of the ego-motion compensated velocity measurements. Furthermore, the radar-based detection performs significantly worse in the terminal environment and environments with a lot of radar reflections.
@@
Lastly, the results of the CenterPoint model training show that the detection quality of diverse object classes (like ’animal’, ’other vehicle’, or ’bus’) is insufficient, while more homogeneous object classes (like ’traffic sign’, ’traffic cone’, or ’pedestrian’) have a higher detection score. Moreover, the size of the objects or the number of objects in the dataset are no sufficient predictors for the detection quality, as shown by the ’truck’ class. In addition, the results suggest that the detection quality decreases with increasing distance and that rainy or foggy weather conditions negatively affect the model’s performance. The model also performs significantly worse in tunnels or rural areas. However, safe autonomous trucking requires robust perception under all environmental conditions, which is why we release MAN TruckScenes to promote research in this area.
PETR | ||||||
AP | ATE | ASE | AOE | AVE | AAE | |
car | 0.03 | 1.26 | 0.22 | 0.20 | 4.18 | 0.12 |
truck | 0.01 | 1.27 | 0.35 | 0.11 | 2.17 | 0.20 |
bus | 0.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 |
trailer | 0.19 | 0.67 | 0.19 | 0.04 | 1.84 | 0.12 |
other_vehicle | 0.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 |
pedestrian | 0.00 | 1.25 | 0.48 | 1.19 | 0.30 | 0.63 |
motorcycle | 0.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 |
bicycle | 0.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 |
traffic_cone | 0.03 | 1.31 | 0.51 | - | - | - |
barrier | 0.01 | 1.46 | 0.90 | 0.07 | - | - |
animal | 0.00 | 1.00 | 1.00 | 1.00 | 1.00 | - |
traffic_sign | 0.00 | 1.29 | 0.61 | 0.50 | - | 0.02 |
RadarGNN | ||||||
AP | ATE | ASE | AOE | AVE | AAE | |
car | 0.36 | 0.53 | 0.22 | 1.48 | 26.50 | 0.02 |
truck | 0.20 | 0.87 | 0.42 | 1.61 | 20.67 | 0.02 |
bus | 0.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 |
trailer | 0.27 | 0.66 | 0.15 | 1.09 | 18.86 | 0.02 |
other_vehicle | 0.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 |
pedestrian | 0.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 |
motorcycle | 0.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 |
bicycle | 0.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 |
traffic_cone | 0.00 | 1.00 | 1.00 | - | - | - |
barrier | 0.00 | 1.00 | 1.00 | 1.00 | - | - |
animal | 0.00 | 1.00 | 1.00 | 1.00 | 1.00 | - |
traffic_sign | 0.02 | 0.64 | 0.92 | 1.27 | - | 0.08 |
@@
car | 0.33 | 0.29 | 0.15 | 0.07 | 3.95 | 0.17 |
truck | 0.23 | 0.42 | 0.16 | 0.03 | 2.62 | 0.39 |
bus | 0.06 | 0.49 | 0.11 | 0.18 | 4.28 | 0.01 |
trailer | 0.34 | 0.42 | 0.15 | 0.03 | 3.74 | 0.37 |
other_vehicle | 0.14 | 0.65 | 0.34 | 0.15 | 0.85 | 0.32 |
pedestrian | 0.45 | 0.29 | 0.39 | 0.81 | 0.43 | 0.37 |
motorcycle | 0.22 | 0.33 | 0.22 | 0.08 | 5.57 | 0.03 |
bicycle | 0.06 | 0.27 | 0.27 | 0.46 | 2.13 | 0.13 |
traffic_cone | 0.55 | 0.28 | 0.37 | - | - | - |
barrier | 0.42 | 0.19 | 0.62 | 0.03 | - | - |
@@
traffic_sign | 0.42 | 0.28 | 0.44 | 0.21 | - | 0.03 |
Modality | - | - | - | - | |
PETR | camera | 0.06 / 0.13 | 0.04 / 0.13 | 0.03 / 0.12 | 0.02 / 0.12 |
RadarGNN | radar | 0.09 / 0.12 | 0.08 / 0.12 | 0.08 / 0.11 | 0.07 / 0.11 |
CenterPoint | lidar | 0.48 / 0.52 | 0.40 / 0.48 | 0.30 / 0.43 | 0.27 / 0.41 |
PETR | RadarGNN | CenterPoint | ||
modality | camera | radar | lidar | |
weather | clear | 0.02 / 0.11 | 0.05 / 0.10 | 0.30 / 0.42 |
overcast | 0.02 / 0.11 | 0.09 / 0.12 | 0.19 / 0.36 | |
fog | 0.05 / 0.13 | 0.15 / 0.15 | 0.15 / 0.20 | |
other_weather | - / - | - / - | - / - | |
rain | 0.03 / 0.11 | 0.09 / 0.12 | 0.16 / 0.23 | |
snow | - / - | - / - | - / - | |
hail | - / - | - / - | - / - | |
area | terminal | 0.02 / 0.09 | 0.00 / 0.04 | 0.30 / 0.32 |
parking | - / - | - / - | - / - | |
highway | 0.02 / 0.11 | 0.09 / 0.12 | 0.19 / 0.31 | |
city | 0.02 / 0.07 | 0.02 / 0.06 | 0.18 / 0.29 | |
residential | 0.01 / 0.06 | 0.03 / 0.07 | 0.12 / 0.28 | |
rural | 0.03 / 0.10 | 0.07 / 0.11 | 0.06 / 0.15 | |
other_area | - / - | - / - | - / - | |
daytime | morning | 0.02 / 0.11 | 0.08 / 0.11 | 0.17 / 0.36 |
noon | 0.02 / 0.12 | 0.06 / 0.11 | 0.33 / 0.44 | |
evening | - / - | - / - | - / - | |
night | 0.02 / 0.10 | 0.06 / 0.09 | 0.17 / 0.24 | |
season | spring | - / - | - / - | - / - |
summer | 0.02 / 0.12 | 0.07 / 0.11 | 0.34 / 0.42 | |
autumn | 0.02 / 0.11 | 0.08 / 0.11 | 0.20 / 0.35 | |
winter | 0.03 / 0.10 | 0.05 / 0.10 | 0.12 / 0.26 | |
lighting | illuminated | 0.02 / 0.12 | 0.07 / 0.11 | 0.26 / 0.41 |
twilight | 0.05 / 0.09 | 0.06 / 0.10 | 0.19 / 0.27 | |
dark | 0.01 / 0.09 | 0.02 / 0.09 | 0.18 / 0.24 | |
glare | 0.02 / 0.11 | 0.09 / 0.12 | 0.16 / 0.23 | |
other_lighting | - / - | - / - | - / - | |
structure | regular | 0.02 / 0.12 | 0.06 / 0.10 | 0.29 / 0.39 |
underpass | 0.04 / 0.12 | 0.11 / 0.13 | 0.13 / 0.28 | |
bridge | 0.03 / 0.11 | 0.09 / 0.12 | 0.14 / 0.30 | |
overpass | 0.03 / 0.12 | 0.09 / 0.12 | 0.24 / 0.31 | |
tunnel | 0.03 / 0.08 | 0.07 / 0.09 | 0.12 / 0.16 | |
const | unchanged | 0.02 / 0.12 | 0.07 / 0.11 | 0.25 / 0.40 |
roadworks | 0.04 / 0.12 | 0.10 / 0.11 | 0.23 / 0.28 |